# Physical aspects of symmetry breaking in an ideal Bose gas at thermal equilibrium

### v2

@@ Abstract
The theory of ideal gases is supplemented by a numerical quantum field description with a two-dimensional non-local order parameter that allows the modeling of wave-like atomic correlations and interference effects in the limit of low atomic densities. From the present model, it is possible to explain certain new fundamental symmetry aspects of ideal and very weakly interacting Bose gases, like the forward propagation of time and the relation to the breaking and preservation of phase gauge symmetry in solids. In the present formalism, the propagation of one-directional time arises from the pre-defined and equivalent convergence of independent quantum fields towards the Boltzmann equilibrium, and it is shown that Glauber coherent states are related to the definition of the quantized field. Coherently coupling condensate and non-condensate parts as a direct consequence of the increasing quantum coherence time between the different quantum field components in the Bose gas from cooling to below the critical temperature, the present model describes symmetry breaking, which is originally known from the definition of a specific gauge field from Elitzur’s theorem for local gauge fields, as a global physical rather than a purely formal mathematical process.
@@ I Introduction
The theory of ideal gases shows a long history of modeling aggregate states and phase transitions of many-particle quantum systems confined to external trapping potentials such as photonic Bose-Einstein condensates ref-1 ; ref-2 . Results and concepts of ideal gas theory have, in particular, been extensively applied and reformulated to address fundamental questions of condensed matter theory with particles of definite masses, such as the condensation of atoms to a Bose-Einstein condensate below the critical temperature ref-3 . Though the assumptions of kinetic ideal gas modeling make it possible to understand the characteristics and dynamics of a very weakly interacting many-particle atomic quantum system, finding a parameter regime of low enough densities and large temperatures that still shows condensation of atoms into a macroscopic ground state population of the typically energetically lowest-lying single-particle energy mode in the particle spectrum in a real experiment often makes it challenging to produce reliable quantitative results for interacting systems when only calculating numerical results for non-interacting particles ref-4 ; ref-5 .
@@
Theoretically, the theory of ideal gases provides a straightforward tool to understand and model the dynamics of many particles for the entire range below and above the critical temperature for Bose-Einstein condensation ref-6 ; ref-7 . Such experimental realizations of Bose-Einstein condensates with minimal interaction strength (described within first-order perturbation theory by the s-wave scattering length) are typically well-covered by the model results of kinetic gas theories, i.e. in the limit of low densities and high temperatures at zero interaction strength ref-8 . Yet the fundamental results and applicational practices of ideal gas theory are well established, so far, more needs to be known about simple theoretical formulations to explain the relation of reversible microscopic many-particle dynamics and the time direction of many-body quantum systems ref-9 ; ref-10 .
@@
In the context of ideal gases, irreversibility is understood as entropy maximization at the Boltzmann equilibrium, as formulated by the second law of thermodynamics. A daily observation of experiments is that time passes irreversibly in one direction, which can e. g. be counted with an atomic clock that measures the pings of oscillating photons or atoms, respectively, at the edges of a microwave cavity or by the number of decays of atoms in an atomic crystal, to define a unit of time. However, since non-interacting gases do not obey dissipation or inelastic collisions, the dynamics of ideal many-body processes are completely reversible in time, and the fundamental process for breaking time or phase gauge symmetries must not depend on the interactions of a many-particle system ref-11 .
It is, therefore, in particular, not trivial to combine the theory of ideal gases with the concept of forward propagating time - since, as argued before, time principally obeys a (reflection) symmetry for non-interacting atomic systems. Relating ideal gases to the concept of spontaneous symmetry breaking in Bose-Einstein condensates in terms of breaking the phase gauge invariance in the quantum system is another delicate task ref-12 , and one may further ask whether or if the forward propagation of a time variable as defined in physical units is related to a sort of broken phase gauge symmetry in a non-interacting many-body quantum system? What is the relation of one-directional time propagation, symmetry breaking at the phase transition for Bose-Einstein condensation, and the occurrence of a quantum field with a well-defined, but presumably random phase that describes the wave nature of the quantum system?
@@
To answer these questions that are currently not straightforwardly explained by the standard theory of ideal gases, the present work explains the context of the fundamental results of a number-conserving quantum field theory to representations of Bose-Einstein condensates in terms of Glauber coherent states. First, the symmetry preserving unitary time evolution according to the ideal gas theory is introduced, before extensions to the description of an ideal gas in terms of a non-local order parameter in complex number representation is explained. In the framework of this non-local quantum field theory for a total number of $N$ particles at a definite temperature $T$, decay, and oscillation rates associated with partial chemical potentials can be numerically derived in complex number representation. This quantum model, in particular, explains the one-directional propagation of time with an average zero phase of the quantum wave field from the interference of ideally disjoint forward and backward propagating wave fields at the Boltzmann equilibrium ref-9 with vanishing total chemical potential. The scaling of the coherence time indicates that the total quantum field is composed of a joint combination of two field components (a field generated from the condensate mode population and a second field created from the rest of the single particle modes) leading to a symmetric total field distribution that builds an extended ring in the complex plane above the critical temperature. This field decays into a point-like field of zero average value as a consequence of the increasing coherence time between condensate and non-condensate subspaces below the critical temperature when the coherence time approaches values of the minimal time uncertainty.
@@
The presented theory allows the numerical calculation for realizations of the quantum field that defines Glauber coherent states for an ideal gas below as well as above the critical temperature for Bose-Einstein condensation ref-13 . Thereby, it is shown ab initio that the phase of the distribution of coherent field states is distributed around zero phases at the Boltzmann equilibrium with a constant total number of $N$ particles at a definite temperature $T$ and vanishing total chemical potential $\mu$, while phase-coherent states with arbitrary phase can, in principle, be realized from Monte-Carlo simulations of the quantum field that follows equilibrium states with lower total entropy corresponding to non-zero values of the chemical potential $\mu\neq 0$.
@@
Finally, this explains that forward time propagation and spontaneously broken gauge symmetries of the Schrödinger equation’s gauge fields occur as a physical process below the critical temperature, originating from quantum coherence of the quantum states in a theory with non-local order parameters that gauges the average total field to zero and further implements the physical boundary condition of a Boltzmann thermal equilibrium into the theoretical model.
@@ II Theoretical Model
In the following, we consider a bosonic gas of $N$ non-interacting atoms confined in an external harmonic trapping potential. A supplementary model for the ideal gas theory is presented that covers the description of the non-reversible forward-time evolution of a non-interacting many-particle system below and above the critical temperature for Bose-Einstein condensation. The model describes the quantum field of an ideal gas in terms of a two-dimensional vector that includes either one order parameter for the quantum system’s forward and backward propagating unitary time evolution. Applying this mathematical structure, interference of the two field components defines the forward propagation of time provided that the system is considered to follow the Boltzmann equilibrium. Time is measured in quantized physical units of seconds, which means that the extended ideal gas model can be interpreted as a model description for elementary quantum clocks. Modeling the quantum field at finite temperature assuming thermal equilibrium leads to numerical realizations of Glauber coherent states with an initial phase defined by the (pre-defined) zero phase of the quantum field.
@@ II.1 Unitary time evolution
with $\omega_{x}$, $\omega_{y}$, and $\omega_{z}$, the trapping frequencies of the external confinement - for very weakly interacting Bose gases. Indeed, given Eqs. (1) - (3), the reversible time evolution of the quantum system can be well described by the Schrödinger equation that describes the time evolution of a many-body quantum state $|\psi_{\bf{k}}\rangle=|\psi_{\bf{k}_{1}}\rangle\otimes...|\psi_{\bf{k}_{N}}\rangle$.
@@ II.2 Quantum field and order parameter
In a gas of $N$ non-interacting atoms, there is ad hoc no theoretical constraint that defines a clear direction of time. The system passes a symmetric unitary time evolution according to the Schrödinger equation. Therefore, the order parameter may be defined for both directions in time, so that the vector
@@
describes an extended point of view for the order parameter, i. e. an order parameter $\psi_{+}(t)=\psi(t)$ that describes the forward and $\psi_{-}(t)=\psi(-t)$ the backward time evolution of the average quantum field. At the thermal Boltzmann equilibrium, which is formally defined by the variable transformation $it=\beta\hbar=-it$ in the formal limit, where $\hbar\rightarrow 0$, the two components of the order parameter effectively equal and the order parameter becomes
@@ II.3 Elementary model for time evolution
An elementary understanding of the time evolution for the Bose gas that is mathematically described in terms of a non-local order parameter can be obtained considering the Bose gas as a very simple model for an elementary quantum clock. From forward and backward propagating photons in a microwave cavity, measurements of the number of outcoupled photons at the edges of the microwave cavity after a certain (oscillation) time interval $\tau_{0}$ from the interaction with the external environment can be obtained. At each measurement cycle, the photon field can hence be assumed to get in thermal equilibrium with a second source from the contact with the external environment, realized by the emission of photons that are outcoupled from the microwave cavity after each oscillation cycle. In such a setup, both the forward as well as the backward propagating field define a (forward and backward) direction of time, and the symmetry-preserving time variable evolves in units of a quantized variable $\tau_{0}$.
It is only at the Boltzmann equilibrium with vanishing chemical potential (which is derived for the equilibrium with the thermal external environment from particle number conservation at definite temperature), where both fields are exactly equal at multiples of this time interval and the total system interacts with the thermal environment. This can be regarded as a measurement process that defines a macroscopic variable with the temperature $T$. Thus, the two field components $\psi_{+}(t)$ and $\psi_{-}(t)$ entail a symmetry property concerning a period $\tau_{0}$ that leads to the (parameter dependent) quantization of the time variable, $t=m\tau_{0}$, with integer values of $m$, where
@@ II.4 Representations with Glauber coherent states
which in the present formalism can be related to the quantum field of the ideal Bose gas at thermal equilibrium with the representation of the field in the complex number plane $\alpha=\sqrt{N}\psi=\sqrt{N}|\psi|{\rm e}^{i\phi}$ and the distribution $P(\alpha)$ which leads to
@@
defines a stationary state of the Bose gas at the Boltzmann equilibrium, given the initial condition $|\psi_{0}\rangle$.
@@ II.5 Correlations of the relative phase
As illustrated in terms of an analytical formulation in the previous chapter, correlations between phases of the forward and backward propagating components of the non-local order parameter give rise to relationships between irreversible macroscopic quantities such as temperature or time with the microscopic description of the Bose gas. Numerically, phase correlations between the two different components $\psi_{+}$ and $\psi_{-}$ can be calculated ab initio from the equation
@@ II.6 Spontaneous phase gauge symmetry breaking
The concept of spontaneous phase gauge symmetry-breaking explains a certain sudden change in the symmetry properties of the gauge fields that describe a solid below the critical temperature. As an example, one may consider a magnet that changes its magnetization from average zero to a well-defined direction with components of spin up and spin down in the z-direction of the magnet below a specific Curie temperature for a characteristic phase transition. While the experimental observation of symmetry breaking is understood straightforwardly, the formulation of a consistent theory that explains spontaneous symmetry breaking as a physical rather than a purely mathematical process is highly non-trivial.
@@
Any initial state of the formal solution in Eq. (13) with a specific initial phase defined by $\psi_{0}=|\psi|{\rm e}^{i\phi_{0}}$ defines a separate and non-unique solution of the Schrödinger equation. The corresponding gauge fields are symmetric as long as no physical constraint is assumed for the quantum field’s chemical potential and are defined by Glauber coherent states that all preserve the symmetry of the Schrödinger equation. Considering only manifolds of the Boltzmann equilibrium with vanishing chemical potential $\mu=0$ leads to projections onto subspaces of the gauge fields as a consequence of a physical constraint, the Boltzmann equilibrium, i.e., to the distribution $P(\alpha)=\delta(\alpha-\alpha_{0})$, without the requirement of formally defining a certain symmetry-breaking gauge for the wave field. Please note, however, that this does not define a symmetry-breaking process of the quantum system, since the symmetry of the total quantum field is preserved. The projection of the total wave field onto the manifold of the Boltzmann equilibrium finally implies the propagation of a one-directional time, i. e. ”the forward direction of time”. From Eqs. (6) - (8), it is thus likely to understand that physical time passes in only one direction, since at the Boltzmann equilibrium, where disjoint forward and backward propagating components of the quantum field are effectively mixed and equal, the partial waves interfere and lead to a well-defined mathematical form of the quantum field that doesn’t diverge numerically also below the critical temperature for Bose-Einstein condensation.
@@
The breaking of the relative phase gauge symmetry between condensate and non-condensate fields during the phase transition can be theoretically understood as an effect of quantum coherence below the critical temperature that changes the total field from a state of the form $\psi=\psi_{0}\cup\psi_{\perp}$ to a coherent superposition of the components $\psi_{0}$ and $\psi_{\perp}$. Considering a quantum field without local representations of the particle’s wave functions facilitates the formulation of the relative symmetry breaking between the different components (aggregate phases) of the quantum system. As will be illustrated numerically ab initio in the following chapter III, in mathematical terms, symmetry breaking in ideal Bose gases below the critical temperature can be understood by the introduction of a non-local order parameter as defined by a coherent sum of condensate and non-condensate fields,
@@
within the framework of the present number-conserving quantum field theory. Since a total number of $N$ non-interacting atoms is assumed, the total average field $\psi=0$ (the gauge of the order parameter of the total field) vanishes (see e.g. Ref. ref-12 ). This is, in particular, described and accounted for by Eq. (9) which represents the conservation of the total number of particles. From this assumption, one is able to recognize that the condensate and non-condensate aggregate phases of the Bose gas are separated by an angle of $\phi=\pi$, i.e.

### v4

# Physical aspects of symmetry breaking in Bose gases at thermal equilibrium

@@ Abstract
The theory of non-interacting Bose gases is supplemented by a numerical quantum field description with a two-dimensional non-local order parameter that allows the modeling of wave-like atomic correlations and interference effects in the limit of low atomic densities. From the present model, it is possible to explain symmetry aspects of non-interacting and very weakly interacting Bose gases in the limit of fluctuating particle numbers, like the forward propagation of time and the relation to the breaking and preservation of phase gauge symmetry in solids. In the present formalism, the propagation of one-directional time arises from the pre-defined and equivalent convergence of independent quantum fields towards the Boltzmann equilibrium, and it is shown that Glauber coherent states are related to the definition of the quantized field. Coherently coupling condensate and non-condensate parts as a direct consequence of the increasing quantum coherence time between the different quantum field components in the Bose gas from cooling to below the critical temperature, the present model describes symmetry breaking, which is originally known from the definition of a specific gauge field from Elitzur’s theorem for local gauge fields, as a global physical rather than a purely formal mathematical process.
@@ I Introduction
The theory of non-interacting Bose gases shows a long history of modeling aggregate states and phase transitions of many-particle quantum systems confined to external trapping potentials such as photonic Bose-Einstein condensates ref-1 ; ref-2 . For system parameters that assume a constant number of particles in the Bose gas, results and concepts of ideal gas theory in the canonical ensemble have, in particular, been extensively applied and reformulated to address fundamental questions of condensed matter theory with particles of definite masses, such as the condensation of atoms to a Bose-Einstein condensate below the critical temperature ref-3 . Though the assumptions of kinetic ideal gas modeling make it possible to understand the characteristics and dynamics of a very weakly interacting many-particle atomic quantum system, finding a parameter regime of low enough densities and large temperatures that still shows condensation of atoms into a macroscopic ground state population of the typically energetically lowest-lying single-particle energy mode in the particle spectrum in a real experiment often makes it challenging to produce reliable quantitative results for interacting systems when only calculating numerical results for non-interacting particles ref-4 ; ref-5 .
@@
Theoretically, the theory of ideal gases in the canonical ensemble provides a straightforward tool to understand and model the dynamics of many particles for the entire range below and above the critical temperature for Bose-Einstein condensation ref-6 ; ref-7 . Such experimental realizations of Bose-Einstein condensates with minimal interaction strength (described within first-order perturbation theory by the s-wave scattering length) are typically well-covered by the model results of kinetic gas theories, i.e. in the limit of low densities and high temperatures at zero interaction strength ref-8 . Yet the fundamental results and applicational practices of ideal gas theory are well established, so far, more needs to be known about simple theoretical formulations to explain the relation of reversible microscopic many-particle dynamics and the time direction of many-body quantum systems ref-9 ; ref-10 .
@@
In the context of non-interacting gases, irreversibility is understood as entropy maximization at the Boltzmann equilibrium, as formulated by the second law of thermodynamics. A daily observation of experiments is that time passes irreversibly in one direction, which can e. g. be counted with an atomic clock that measures the pings of oscillating photons or atoms, respectively, at the edges of a microwave cavity or by the number of decays of atoms in an atomic crystal, to define a unit of time. However, since non-interacting gases do not obey dissipation or inelastic collisions, the dynamics of ideal many-body processes are completely reversible in time, and the fundamental process for breaking time or phase gauge symmetries must not depend on the interactions of a many-particle system ref-11 .
For that purpose, however, the theory of ideal gases that assumes a constant number of particles in the Bose gas has to be extended to the case of lowly fluctuating particle numbers. In this limit, it is in particular possible to combine the theory of ideal gases with the concept of forward propagating time, which is a highly non-trivial conceptual task, since, as argued before, time principally obeys a (reflection) symmetry for non-interacting atomic systems. Relating the theory of non-interacting Bose gases to the concept of spontaneous symmetry breaking in Bose-Einstein condensates in terms of breaking the phase gauge invariance in the quantum system ref-12 requires the assumption of a fluctuating total number of particles in different realizations of the physical setup in a temperature limit below the critical temperature, where the canonical and the grand-canonical ensembles are no longer equivalent. To supplement physical aspects of spontaneously broken gauge symmetries in non-interacting Boses gases that are currently not straightforwardly explained by the standard theory of ideal gases, the present work explains the context of the fundamental results of a number-conserving quantum field theory to representations of Bose-Einstein condensates in terms of Glauber coherent states in the limit, where the particle number is finite and allowed to fluctuate slightly around the mean value defined by a macrcoscopic average value.
@@
In the sequel to the present theory, first, the symmetry-preserving unitary time evolution according to the theory of non-interacting Bose gases is introduced, before extensions to the description of non-interacting Bose gases in terms of a non-local order parameter in complex number representation are explained. In the framework of this non-local quantum field theory for a total average number of $N$ particles at a definite temperature $T$, decay, and oscillation rates associated with partial chemical potentials can be numerically derived in complex number representation. This quantum model, in particular, explains the one-directional propagation of time with an average zero phase of the quantum wave field from the interference of ideally disjoint forward and backward propagating wave fields at the Boltzmann equilibrium ref-9 with vanishing total chemical potential. The scaling of the coherence time indicates that the total quantum field is composed of a joint combination of two field components (a field generated from the condensate mode population and a second field created from the rest of the single particle modes), leading to a symmetric total field distribution that builds an extended ring in the complex plane above the critical temperature. This field decays into a point-like field of zero average value as a consequence of the increasing coherence time between condensate and non-condensate subspaces below the critical temperature, when the coherence time approaches values of the minimal time uncertainty.
@@
The presented theory allows the numerical calculation for realizations of the quantum field that defines Glauber coherent states for a non-interacting Bose gas below as well as above the critical temperature for Bose-Einstein condensation ref-13 . Thereby, it is shown ab initio that the phase of the distribution of coherent field states is distributed around zero phases at the Boltzmann equilibrium with a constant total number of $N$ particles at a definite temperature $T$ and vanishing total chemical potential $\mu$, while phase-coherent states with arbitrary phase can, in principle, be realized from Monte-Carlo simulations of the quantum field that follows equilibrium states with lower total entropy corresponding to non-zero values of the chemical potential $\mu\neq 0$.
@@
Finally, this explains that forward time propagation and spontaneously broken gauge symmetries of the Schrödinger equation’s gauge fields occur as a physical process below the critical temperature, originating from quantum coherence of the quantum states in a theory with non-local order parameters that gauges the average total field to zero and further implements the physical boundary condition of a Boltzmann thermal equilibrium into the theoretical model. The model applies in the limit of non-interacting Bose gases with also lowly fluctuating particle numbers, i.e. nearly ideal Bose gases below the critical temperature.
@@ II Theoretical Model
In the following, we consider a bosonic gas of $N$ non-interacting atoms confined in an external harmonic trapping potential. A supplementary model for the ideal gas theory is presented that covers the description of the non-reversible forward-time evolution of a non-interacting many-particle system below and above the critical temperature for Bose-Einstein condensation in the limit, where the total particle number is allowed to fluctuate from one realization to the next.
The model describes the quantum field of a non-interacting Bose gas in terms of a two-dimensional vector that includes either one order parameter for the quantum system’s forward and backward propagating unitary time evolution. Applying this mathematical structure, the interference of the two field components defines the forward propagation of time provided that the system is considered to follow the Boltzmann equilibrium. Time is measured in quantized physical units of seconds, which means that the extended ideal gas model can be interpreted as a model description for elementary quantum clocks. Modeling the quantum field at finite temperature, assuming thermal equilibrium, leads to numerical realizations of Glauber coherent states with an initial phase defined by the (pre-defined) zero phase of the quantum field.
@@ II.1 Unitary time evolution
with $\omega_{x}$, $\omega_{y}$, and $\omega_{z}$, the trapping frequencies of the external confinement, for very weakly interacting Bose gases. Indeed, given Eqs. (1) - (3), the reversible time evolution of the quantum system can be well described by the Schrödinger equation that describes the time evolution of a many-body quantum state $|\psi_{\bf{k}}\rangle=|\psi_{\bf{k}_{1}}\rangle\otimes...|\psi_{\bf{k}_{N}}\rangle$.
@@ II.2 Quantum field and order parameter
In a gas of $N$ non-interacting atoms, there is no ad hoc theoretical constraint that defines a clear direction of time. The system passes a symmetric unitary time evolution according to the Schrödinger equation. Therefore, the order parameter may be defined for both directions in time, so that the vector
@@
describes an extended point of view for the order parameter, i. e. an order parameter $\psi_{+}(t)=\psi(t)$ that describes the forward and $\psi_{-}(t)=\psi(-t)$ the backward time evolution of the average quantum field. At the thermal Boltzmann equilibrium, which is formally defined by the variable transformation $it=\beta\hbar=-it$ in the formal limit, where $\hbar\rightarrow 0$, the two components of the order parameter effectively equal, and the order parameter becomes
@@ II.3 Elementary model for time evolution
An elementary understanding of the time evolution for the Bose gas that is mathematically described in terms of a non-local order parameter can be obtained by considering the Bose gas as a very simple model for an elementary quantum clock. From forward and backward propagating photons in a microwave cavity, measurements of the number of outcoupled photons at the edges of the microwave cavity after a certain (oscillation) time interval $\tau_{0}$ from the interaction with the external environment can be obtained. At each measurement cycle, the photon field can hence be assumed to get in thermal equilibrium with a second source from the contact with the external environment, realized by the emission of photons that are outcoupled from the microwave cavity after each oscillation cycle. In such a setup, both the forward as well as backward propagating fields define a (forward and backward) direction of time, and the symmetry-preserving time variable evolves in units of a quantized variable $\tau_{0}$.
It is only at the Boltzmann equilibrium with vanishing chemical potential (which is derived for the equilibrium with the thermal external environment from particle number conservation at a definite temperature), where both fields are exactly equal at multiples of this time interval, and the total system interacts with the thermal environment. This can be regarded as a measurement process that defines a macroscopic variable with the temperature $T$. Thus, the two field components $\psi_{+}(t)$ and $\psi_{-}(t)$ entail a symmetry property concerning a period $\tau_{0}$ that leads to the (parameter-dependent) quantization of the time variable, $t=m\tau_{0}$, with integer values of $m$, where
@@ II.4 Representations with Glauber coherent states
which in the present formalism can be related to the quantum field of the non-interacting Bose gas at thermal equilibrium with the representation of the field in the complex number plane $\alpha=\sqrt{N}\psi=\sqrt{N}|\psi|{\rm e}^{i\phi}$ and the distribution $P(\alpha)$ which leads to
@@
defines a stationary state of the Bose gas, given the initial condition $|\psi_{0}\rangle$.
@@ II.5 Correlations of the relative phase
As illustrated in terms of an analytical formulation in the previous chapter, correlations between phases of the forward and backward propagating components of the non-local order parameter gives rise to relationships between irreversible macroscopic quantities, such as temperature or time, with the microscopic description of the Bose gas. Numerically, phase correlations between the two different components $\psi_{+}$ and $\psi_{-}$ can be calculated ab initio from the equation
@@ II.6 Spontaneous phase gauge symmetry breaking
The concept of spontaneous phase gauge symmetry-breaking explains a certain sudden change in the symmetry properties of the gauge fields that describe a solid below the critical temperature. As an example, one may consider a magnet that changes its magnetization from an average zero to a well-defined direction with components of spin up and spin down in the z-direction of the magnet below a specific Curie temperature for a characteristic phase transition. While the experimental observation of symmetry breaking is understood straightforwardly, the formulation of a consistent theory that explains spontaneous symmetry breaking as a physical rather than a purely mathematical process is highly non-trivial.
@@
Any initial state of the formal solution in Eq. (13) with a specific initial phase defined by $\psi_{0}=|\psi|{\rm e}^{i\phi_{0}}$ defines a separate and non-unique solution of the Schrödinger equation. The corresponding gauge fields are symmetric as long as no physical constraint is assumed for the quantum field’s chemical potential and are defined by Glauber coherent states that all preserve the symmetry of the Schrödinger equation. Considering only manifolds of the Boltzmann equilibrium with vanishing chemical potential $\mu=0$ leads to projections onto subspaces of the gauge fields as a consequence of a physical constraint, the Boltzmann equilibrium, i.e., to the distribution $P(\alpha)=\delta(\alpha-\alpha_{0})$, without the requirement of formally defining a certain symmetry-breaking gauge for the wave field. Please note, however, that this does not define a symmetry-breaking process of the quantum system, since the symmetry of the total quantum field is preserved. The projection of the total wave field onto the manifold of the Boltzmann equilibrium finally implies the propagation of a one-directional time, i. e. ”the forward direction of time”. From Eqs. (6) - (8), it is thus likely to understand that physical time passes in only one direction, since at the Boltzmann equilibrium, where disjoint forward and backward propagating components of the quantum field are effectively mixed and equal, the partial waves interfere and lead to a well-defined mathematical form of the quantum field that doesn’t diverge numerically, also below the critical temperature for Bose-Einstein condensation.
@@
The breaking of the relative phase gauge symmetry between condensate and non-condensate fields during the phase transition can be theoretically understood as an effect of quantum coherence below the critical temperature that changes the total field from a state of the form $\psi=\psi_{0}\cup\psi_{\perp}$ to a coherent superposition of the components $\psi_{0}$ and $\psi_{\perp}$. Considering a quantum field without local representations of the particle’s wave functions facilitates the formulation of the relative symmetry breaking between the different components (aggregate phases) of the quantum system. As will be illustrated numerically ab initio in the following chapter III, in mathematical terms, symmetry breaking in non-interacting Bose gases below the critical temperature can be understood by the introduction of a non-local order parameter as defined by a coherent sum of condensate and non-condensate fields,
@@
within the framework of the present number-conserving quantum field theory. Since a total number of $N$ non-interacting atoms is assumed, the total average field $\psi=0$ (the gauge of the order parameter of the total field) vanishes (see e.g. Ref. ref-12 ). This is, in particular, described and accounted for by Eq. (9) which represents the conservation of the average total number of particles. From this assumption, one can recognize that the condensate and non-condensate aggregate phases of the Bose gas are separated by an angle of $\phi=\pi$, i.e.